# HASD: Hierarchical Adaption for pathology Slide-level Domain-shift

### v1

@@ Abstract
Domain shift is a critical problem for pathology AI as pathology data is heavily influenced by center-specific conditions. Current pathology domain adaptation methods focus on image patches rather than WSI, thus failing to capture global WSI features required in typical clinical scenarios. In this work, we address the challenges of slide-level domain shift by proposing a Hierarchical Adaptation framework for Slide-level Domain-shift (HASD). HASD achieves multi-scale feature consistency and computationally efficient slide-level domain adaptation through two key components: (1) a hierarchical adaptation framework that integrates a Domain-level Alignment Solver for feature alignment, a Slide-level Geometric Invariance Regularization to preserve the morphological structure, and a Patch-level Attention Consistency Regularization to maintain local critical diagnostic cues; and (2) a prototype selection mechanism that reduces computational overhead. We validate our method on two slide-level tasks across five datasets, achieving a 4.1% AUROC improvement in a Breast Cancer HER2 Grading cohort and a 3.9% C-index gain in a UCEC survival prediction cohort. Our method provides a practical and reliable slide-level domain adaption solution for pathology institutions, minimizing both computational and annotation costs.
@@ 1 Introduction
Digital pathology has experienced significant attention in recent years, driven by the widespread adoption of whole-slide imaging (WSI) scanners and rapid progress in artificial intelligence (AI) research [15]. This trend has facilitated notable advancements in many AI-based pathology tasks [25, 6]. However, a major challenge in pathology is domain shift, arising from differences in pathology centers, including imaging variances (varying staining protocols, digitization procedures), label discrepancies (patient populations), and other factors. Domain shift can downgrade model performances when transferring a model to a different clinical site, potentially compromising patient safety [11, 13].
@@ 2.1.3 Optimal Transport
Traditional OT is not well-suited for pathology slide-level domain adaptation, as it incurs significant computational overhead [24] and treats each patch as an isolated point, thereby ignoring the structured slide morphology and failing to preserve multi-scale feature consistency [14]. In the following section, we introduce our method to address these challenges.
@@ 3 Experiments and Results
The prediction of the HER2 status in breast cancer from hematoxylin & eosin (H&E) stained tumor tissue is a challenging slide-level task [12]. We evaluate our method on multi-institutional breast cancer slides from three centers: Yale Hospital (Yale, n=192) [7], The Cancer Genome Atlas (TCGA, n=182) [23], and a local hospital (INT, n=77).
@@ 3.1 Quantitative Evaluation
Her2 Grading Task Methods ID Average OOD Average MI-SimpleShot 81.6 70.8(10.8$\downarrow$) SCDA 82.5 72.0(10.5$\downarrow$) ABMIL 85.8 75.8(10.0$\downarrow$) ReinhardNorm 85.3 75.3(10.0$\downarrow$) Ours 86.1 79.9(6.2$\downarrow$) Survival Prediction Task Methods ID Average OOD Average MI-SimpleShot 64.9 52.9(12.0$\downarrow$) SCDA 65.3 54.7(10.6$\downarrow$) ABMIL 68.1 57.5(10.6$\downarrow$) ReinhardNorm 68.5 56.3(12.2$\downarrow$) Ours 68.1 61.4(6.7$\downarrow$)
@@
Evaluating the ID vs. OOD gap is essential for assessing domain shift robustness. A smaller gap indicates better generalization. Table 3.1 summarizes the differences. For HER2 Grading, MI-SimpleShot suffers the largest drop ($10.8\%$ AUC), while our method maintains the smallest gap ($6.2\%$ AUC) with the highest OOD score, demonstrating strong generalization. In Survival Prediction, domain shift is further influenced by label prevalence differences. All other methods drops $>10\%$ CI, while our method, aided by partial DAS, achieves the best OOD performance with a smaller gap ($6.7\%$ CI), proving its effectiveness.
@@ 3.2 Ablation study
DAS # Prototype $k$ PACR SGIR Grading Average Survial Prediction Average $\checkmark$ 1 75.4 58.2 $\checkmark$ 5 76.2(0.8$\uparrow$) 58.6(0.4$\uparrow$) $\checkmark$ 10 76.1(0.1$\downarrow$) 59.1(0.5$\uparrow$) $\checkmark$ 10 $\checkmark$ 79.6(3.5$\uparrow$) 58.6(0.5$\downarrow$) $\checkmark$ 10 $\checkmark$ $\checkmark$ 79.9(0.3$\uparrow$) 60.1(1.5$\uparrow$) Partial DAS 10 $\checkmark$ $\checkmark$ - 61.4(1.3$\uparrow$)
@@ 4 Conclusion
Domain shift is a critical problem for pathology AI as pathology data is heavily influenced by center-specific conditions. Current pathology domain adaptation methods focus on image patches rather than WSI, thus failing to capture global WSI features required in typical clinical scenarios. We, for the first time, introduced a slide-level domain adaption method HASD, which incorporates a hierarchical adaptation design to maintain multi-scale feature consistency and a prototype selection mechanism to achieve computationally efficiency. Our method outperformed conventional DA methods in two tasks (Her2 grading and Survival) by 4.1% and 3.9%, when models were trained and tested in different centers, while preserving in-domain performance. We believe that our research on slide-level domain adaptation will improve the generalizability of future clinical AI models. Code is available at anonymous.

### v2

@@ Abstract
Domain shift is a critical problem for artificial intelligence (AI) in pathology as it is heavily influenced by center-specific conditions. Current pathology domain adaptation methods focus on image patches rather than whole-slide images (WSI), thus failing to capture global WSI features required in typical clinical scenarios. In this work, we address the challenges of slide-level domain shift by proposing a Hierarchical Adaptation framework for Slide-level Domain-shift (HASD). HASD achieves multi-scale feature consistency and computationally efficient slide-level domain adaptation through two key components: (1) a hierarchical adaptation framework that integrates a Domain-level Alignment Solver for feature alignment, a Slide-level Geometric Invariance Regularization to preserve the morphological structure, and a Patch-level Attention Consistency Regularization to maintain local critical diagnostic cues; and (2) a prototype selection mechanism that reduces computational overhead. We validate our method on two slide-level tasks across five datasets, achieving a 4.1% AUROC improvement in a Breast Cancer HER2 Grading cohort and a 3.9% C-index gain in a UCEC survival prediction cohort. Our method provides a practical and reliable slide-level domain adaption solution for pathology institutions, minimizing both computational and annotation costs. Code is available at https://github.com/TumVink/HASD.
@@ 1 Introduction
Digital pathology has experienced significant attention in recent years, driven by the widespread adoption of whole-slide imaging (WSI) scanners and rapid progress in artificial intelligence (AI) research [15]. This trend has facilitated notable advancements in many AI-based pathology tasks [28, 6]. However, a major challenge in pathology is domain shift, arising from differences in pathology centers, including imaging variances (varying staining protocols, digitization procedures), label discrepancies (patient populations), and other factors. Domain shift can downgrade model performances when transferring a model to a different clinical site, potentially compromising patient safety [11, 13, 24, 26].
@@ 2.1.3 Optimal Transport
Traditional OT is not well-suited for pathology slide-level domain adaptation, as it incurs significant computational overhead [27, 25] and treats each patch as an isolated point, thereby ignoring the structured slide morphology and failing to preserve multi-scale feature consistency [14]. In the following section, we introduce our method to address these challenges.
@@ 3 Experiments and Results
The prediction of the HER2 status in breast cancer from hematoxylin & eosin (H&E) stained tumor tissue is a challenging slide-level task [12]. We use AUROC to evaluate our method on multi-institutional breast cancer slides from three centers: Yale Hospital (Yale, n=192) [7], The Cancer Genome Atlas (TCGA, n=182) [23], and Technical University Munich hospital (TUM, n=77).
@@ 3.1 Quantitative Evaluation
[…]
@@
Evaluating the ID vs. OOD gap is essential for assessing domain shift robustness. A smaller gap indicates better generalization. Table 3.1 summarizes the differences. For HER2 Grading, MI-SimpleShot suffers the largest drop ($10.8\%$ AUC), while our method maintains the smallest gap ($6.2\%$ AUC) with the highest OOD score, demonstrating strong generalization. In Survival Prediction, domain shift is further influenced by label prevalence differences. All other methods drops $10\%$ CI, while our method, aided by partial DAS, achieves the best OOD performance with a smaller gap ($6.7\%$ CI), proving its effectiveness.
@@ 3.2 Ablation study
[…]
@@ 4 Conclusion
Domain shift is a critical problem for pathology AI as pathology data is heavily influenced by center-specific conditions. Current pathology domain adaptation methods focus on image patches rather than WSI, thus failing to capture global WSI features required in typical clinical scenarios. We, for the first time, introduced a slide-level domain adaption method HASD, which incorporates a hierarchical adaptation design to maintain multi-scale feature consistency and a prototype selection mechanism to achieve computationally efficiency. Our method outperformed conventional DA methods in two tasks (Her2 grading and Survival) by 4.1% and 3.9%, when models were trained and tested in different centers, while preserving in-domain performance. We believe that our research on slide-level domain adaptation will improve the generalizability of future clinical AI models.